OpenStudy (anonymous):

One to one function h is defined by h(x)=(6x)/(-8+7x). Find h^-1, the inverse of h. Then, give the domain and range of h^-1 using interval notation.

OpenStudy (johnweldon1993):

Do you know how to find the inverse?

OpenStudy (johnweldon1993):

You begin by switching the 'x's and the 'y's So your equation \[y = \frac{ 6x }{ -8 + 7x }\] Will become \[x = \frac{ 6y }{ -8 + 7y }\] Now you just solve for 'y' again...

OpenStudy (johnweldon1993):

So multiply both sides by -8 + 7y to get rid of that fraction on the right hand side \[x(-8 + 7y) = 6y\] Distribute out that 'x' \[-8x + 7xy = 6y\] Now you want to solve for 'y' so move everything with a 'y' to 1 side \[6y - 7xy = -8x\] Factor out a 'y' from the left hand side \[y(6 - 7x) = -8x\] And finally to solve for 'y' divide both sides of the equation by 6 - 7x \[y = \frac{ -8x }{ 6 - 7x }\] That is your inverse....can you tell me the domain or range?

OpenStudy (anonymous):

i dont understand how to find the domain and range

OpenStudy (johnweldon1993):

To find the domain...you find out what values of 'x' will make your denominator = 0 so \[-7x + 6 = 0\] Can you solve that equation for 'x'...?

OpenStudy (anonymous):

x would be 6/7 how do you find the range then?

OpenStudy (johnweldon1993):

Well the range would be all real numbers not equal to that number we just found...the domain...

OpenStudy (anonymous):

how would i write that? -infinity, infinity?